OpenStudy (anonymous):

-[8z-(12z+5)=5+(3z+4)

OpenStudy (anonymous):

\[-8z-(12z+5) = 5+(3z+4)\]

OpenStudy (anonymous):

Lets put it nicely first, and I assume we're solving for z?

OpenStudy (anonymous):

Drop the brackets \[-8z-12z-5=5+3z+\]

OpenStudy (anonymous):

Collect like terms \[-23z = 14\]

OpenStudy (anonymous):

Divide \[z = -14/23\]

OpenStudy (phi):

you are missing a closing square bracket in -[8z-(12z+5)=5+(3z+4) where is the ] ?

OpenStudy (anonymous):

I assumed that was a mistake because I thought the same at first haha.